OpenStudy (anonymous):

On a factory floor, 30 out of every 140 toy robots is defective. What percent of the toy robots are defective? Round your answer to the nearest hundredth

OpenStudy (caozeyuan):

30/140=3/14=1.5/7=1.5*1/7=1.5*0.142857=0.21

OpenStudy (anonymous):

30/140=21.43%